OpenStudy (anonymous):

simplify.. X 2 ________ - ________ (x+2)(x-1) (x-4)(x-1)

OpenStudy (anonymous):

\[\frac{x}{(x+2)(x-1)}-\frac{2}{(x-4)(x-1)}\]\[\frac{x}{(x+2)(x-1)} \times \frac{x-4}{x-4}-\frac{2}{(x-4)(x-1)} \times \frac{x+2}{x+2}\]\[\frac{x(x-4)}{(x+2)(x-1)(x-4)}-\frac{2(x+2)}{(x+2)(x-1)(x-4)}\]\[\frac{x^{2}-4x-2x-4}{(x+2)(x-1)(x-4)}\]